3. A proton = 1.60 x 10-19 C, moves through a uniform magnetic field into the page. The proton, moving at a speed of 2.0 x 10 please show work!


e) If instead of a proton an electron were moving at the same speed and direction in the same magnetic field would the electr part c as well please! im really struggling and on a time limit please help!
Answer #1

F = Bov sino Ozgo 3x10-15 = BY (6816-192x106 B = 3x10-15 To6x10-19 x 2x106 B = 6-9375 x 10-15 B = 93.75T Proton - Tv For elec